OpenStudy (anonymous):

x² + 4x - 7 = 0 solve, and show work.

OpenStudy (anonymous):

Okay, I love those exercises. You want to solve for x1 and x2?

OpenStudy (anonymous):

If yes, there is a standard formula Delta = b^2-4ac. x^2 in this case plays the role of a; 4x plays the role of b; and -7 plays the role of c so by using that standard formula we now just have to replace everything... Delta = 4^2 - 4*1*(-7) => Delta = 16 + 28 => Delta = 44 So now that we know Delta is not less than zero, we know it has x1, x2 solutions... The x1, x2 solutions are also calculate by a standard formula which is -b^2 +- sqrt(delta) everything divided by 2a so we'll have: x1 = -16 + sqrt(44) / 2*1 = -4.6833752096446 :-) x2 = -16 - sqrt(44) / 2* 1 = -11.3166247903554 or something like that! :D